# Unlabeled Data or Pre-trained Model: Rethinking Semi-Supervised Learning and Pretrain-Finetuning

### v2

@@ 1 Introduction
Deep neural networks have demonstrated performance comparable to that of humans in certain supervised learning tasks, such as image classification geirhos2018imagenet ; resnet ; alexnet . However, these impressive results rely heavily on large amounts of labeled training data. In many real-world applications, obtaining such labeled data often demands substantial human and financial resources yang2022survey . To reduce the dependence on labeled data and thereby cut costs, researchers have proposed Semi-Supervised Learning (SSL)—an effective approach that leverages unlabeled data to enhance model performance.
Recently, the Pretrain-Finetuning paradigm has gained significant attention, as models pre-trained on large-scale, readily available web data exhibit a strong ability to rapidly adapt to downstream tasks through fine-tuning. In image classification, for example, the semantic alignment acquired during pre-training enables models—such as CLIP CLIP , a pre-trained vision-language model—to perform zero-shot classification without task-specific supervision. Furthermore, building upon the rich knowledge gained during pre-training, various fine-tuning strategies (e.g., prompt tuning) have been proposed to boost task-specific performance with minimal labeled data maple ; promptsrc ; cocoop ; coop .
@@ 2 Related Work
Semi-Supervised Learning. The goal of Semi-Supervised Learning is to reduce the cost of data annotation by leveraging unlabeled data, under the assumption that both labeled and unlabeled samples belong to a predefined set of known classes. The primary objective is to assign accurate ground-truth labels to unlabeled instances. Existing SSL methods can be broadly categorized into three main approaches: Entropy Minimization 22 ; 8 , Consistency Regularization20 ; 24 ; 29 , and Hybrid or Holistic Methods remixmatch ; mixmatch . Specifically, EntMin 22 promotes confident (i.e., low-entropy) predictions on unlabeled samples by introducing an additional loss term that minimizes prediction uncertainty, regardless of class. Pseudo-Labelinglee2013pseudo assigns pseudo-labels to unlabeled data based on the model’s own predictions throughout training. FixMatch fixmatch combines consistency regularization and confidence-based filtering by enforcing consistency between weakly and strongly augmented samples, while only using high-confidence pseudo-labels for training, resulting in significant performance gains with limited labeled data.
Pretrain-Finetuning. The recent progress in pre-trained models has garnered significant attention, particularly with the emergence of Large Language Models like GPT achiam2023gpt and Vision Language Models like CLIP CLIP . These models, pre-trained on vast amounts of data, exhibit remarkable adaptability to downstream tasks, leading to the widespread adoption of the pretrain-finetuning paradigm. In image classification tasks, VLMs Flamingo ; CLIP ; filip are pre-trained on large-scale text-image pairs, leveraging shared embeddings of textual and visual representations to enable zero-shot classification. While these pre-trained VLMs have captured broad, generalized representations of images and text, efficiently adapting them to downstream tasks remains a significant challenge. Inspired by the success of prompt tuning in NLP he2022hyperprompt ; li2021prefix , many researchers have proposed to adapt VLMs by learning the prompts in end-to-end training. As the pioneering work, CoOp coop for the first time introduces learnable prompts to transfer the task-specific knowledge to VLMs. Building on this idea, MaPLe maple and PromptSRC promptsrc extend prompt tuning to both visual and textual encoders, enabling joint vision-language adaptation. In Few-shot SSL, we select CoOp and PromptSRC as the primary baselines.
@@ 3.2 Pretrain-Fintuning
Pre-trained models typically involve two stages before downstream evaluation (see Figure 1(b)). Taking CLIP CLIP as an example, it employs an image encoder $f$ and a text encoder $g$ to extract feature representations $f(I)$ and $g(T)$ from an input image $I$ and its corresponding text caption $T$, respectively. During pre-training, the model minimizes a contrastive loss over $M$ image-text pairs:
@@
To enable rapid adaptation to downstream tasks, VLMs cocoop ; coop fine-tune a set of learnable prompts $P=[t_{1},t_{2},\dots,t_{b}]$, while keeping both the image and text encoders frozen. Given class name embeddings $[c_{1},c_{2},\dots,c_{N}]$, each class representation is formed by concatenating $P$ with $c_{i}$, and the prediction is computed as:
@@ 3.3 Few-Shot SSL
  • Standard Semi-Supervised Setting fixmatch . In this setting, the labels for both subsets belong to the same known class set $\mathcal{Y}^{l}$. For VLMs, which do not rely on unlabeled data for training, the Few-shots’ SSL framework reduces this setting to a standard few-shot scenario, provided the number of labeled samples remains the same.
  • Open-Set Semi-Supervised Setting os1 . In this setting, the unlabeled dataset $\mathcal{D}_{u}$ may include instances from unseen classes, i.e., $\mathcal{Y}^{l}\subset\mathcal{Y}^{u}$ with $\mathcal{Y}^{\text{new}}=\mathcal{Y}^{u}\setminus\mathcal{Y}^{l}$. Since VLMs do not utilize unlabeled data, the open-set setting in Few-shots’ SSL reduces to the semi-supervised setting.
  • Open-World Semi-Supervised Setting niu2024owmatch . This setting builds upon the open-set setting, aiming to classify both known and new classes. For VLMs, pre-training endows them with zero-shot classification capability, allowing for direct classification of both known and new classes. Therefore, the Open-world semi-supervised setting is analogous to the commonly used Base-to-Novel setting cocoop for VLMs.
  • Long-Tailed Semi-Supervised Setting lt1ts . This setting features a highly imbalanced class distribution, where head classes have many samples and tail classes have few. The imbalance ratio (IR) is defined as the sample count ratio between the largest and smallest classes. Since VLMs use a fixed number of samples per class in few-shot settings, the head and tail classes become consistent due to the same number of samples. Thus, the Few-shot SSL framework reduces this setting to a standard few-shot scenario by matching the number of shots to that of the tail classes, ensuring fair comparison.
@@ 4.1 Baselines Selection
We selected FixMatch fixmatch (standard semi-supervised learning), DWD dwusl (open-set semi-supervised learning), OwMatch niu2024owmatch (open-world semi-supervised learning), and CCL zhou2024continuous (long-tailed semi-supervised learning) as representative methods across different settings. For the pretrain–finetuning paradigm, we consider CLIP CLIP , a pre-trained VLM, along with its prompt tuning variants—CoOp coop , the first to introduce prompt tuning for CLIP, and PromptSRC promptsrc , a state-of-the-art method. Including these methods in our comparison allows for a direct observation of the lower and upper bounds for pre-trained VLMs in semi-supervised settings, leading to a more comprehensive analysis.
@@
UDA uda 68.78 $\pm$ 0.43 (10% labels)
FixMatch fixmatch 71.46 $\pm$ 0.52 (10% labels)
ZSCLIP CLIP 66.70 $\pm$ 0.00 (0 labels)
CoOp coop 73.43 $\pm$ 0.12 (2% labels)
PromptSRC promptsrc 74.03 $\pm$ 0.09 (2% labels)
@@
Datasets and Experimental Setup. We evaluate both SSL and pre-trained VLMs on CIFAR-10/100 cifar , STL-10 stl10 , and ImageNet imagenet , under varying annotation levels and augmentation strategies. For SSL methods (e.g., UDA uda , MixMatch mixmatch ), we report the standard results as presented in FixMatch fixmatch . For VLMs, we simulate few-shot learning by training on the same labeled data only, without accessing unlabeled samples. All experiments on VLMs are conducted using the ViT-B/16 backbone and evaluated over three random seeds. Full implementation details are provided in Appendix A.2.
Experiment Results. As shown in Table 1, SSL methods generally perform well on lower-resolution datasets such as CIFAR-10/100 (32 $\times$ 32), surpassing zero-shot CLIP and its fine-tuning variants. This advantage stems from the difficulty pre-trained VLMs face in extracting meaningful features from lower-resolution inputs. In contrast, pre-trained VLMs excels on higher-resolution datasets like STL-10 (96 $\times$ 96), where even zero-shot CLIP surpasses the best-performing SSL baselines. Moreover, as shown in Table 2(a), on the ImageNet dataset, FixMatch requires a significantly larger proportion of labeled data (around 10%) but still lags behind in accuracy. By comparison, VLMs achieve over 73% accuracy using only 2% of labeled samples, with performance improving as more labeled data becomes available. In addition to accuracy, SSL methods typically involve much longer training times. For instance, training SSL models on CIFAR datasets takes several hours remixmatch ; fixmatch , whereas prompt tuning approaches for VLMs maple ; promptsrc ; coop can be completed in just a few minutes. These observations highlight the complementary strengths of SSL and pretrain-finetuning paradigms, with the latter generally offering a higher performance ceiling and faster training speed.
Analysis. To investigate how input resolution affects pre-trained models’ performance, we evaluate several methods on the STL-10 dataset at varying resolutions (see Table 2(b)). As resolution increases, all VLMs exhibit substantial accuracy gains, highlighting their sensitivity to image resolution. We also experiment with different visual encoder backbones (presented in Appendix B) and find that ViT-based architectures consistently outperform ResNet-50 across all resolutions, indicating that the performance degradation is not due to an intrinsic limitation of Vision Transformer dosovitskiy2020image . In fact, it stems from the fixed input size requirement (e.g., 224 $\times$ 224) of VLMs: low-resolution images must be upsampled via interpolation to satisfy this requirement, a process that inevitably introduces distortions and impairs the models’ ability to extract reliable visual features.
@@
Datasets and Experimental Setup. We evaluate both paradigms on standard open-set benchmarks, including CIFAR-10/100 cifar , ImageNet-30 hendrycks2019using , and ImageNet-100 cao2021open . For CIFAR-10/100, we use 100 labeled images per class from CIFAR-10 as the supervised data and the entire CIFAR-100 dataset as the unsupervised data. For ImageNet-30 and ImageNet-100, we adopt a 50-shot setting, using 50 labeled samples per class from the selected in-distribution (ID) classes. It is important to note that the test set in this setting contains only those classes present in the training data. For the pretrain–finetuning paradigm, the experimental setup and VLM performance remain consistent with those described in the semi-supervised setting. Further details can be found in Appendix A.3.
@@
FixMatch fixmatch 71.5 50.4 49.5 39.6 23.5 20.3 65.8 36.7 34.9
DS3d3l 77.6 45.3 40.2 55.1 23.7 24.0 71.2 32.5 30.8
CGDL cgdl 72.3 44.6 39.7 49.3 22.5 23.6 67.3 33.8 31.9
DTC dtc 53.9 39.5 38.3 31.3 22.9 18.3 25.6 20.8 21.3
RankStats RankStats 86.6 81.0 82.9 36.4 28.4 23.1 47.3 28.7 40.3
SimCLR simclr 58.3 63.4 51.7 28.6 21.1 22.3 39.5 35.7 36.9
ORCA orca 88.2 90.4 89.7 66.9 43.0 48.1 89.1 72.1 77.8
NACH nach 89.5 92.2 91.3 68.7 47.0 52.1 91.0 75.5 79.6
OpenLDN rizve2022openldn 95.7 95.1 95.4 73.5 46.8 60.1 89.6 68.6 79.1
OpenCon sun2022opencon 89.3 91.1 90.4 69.1 47.8 52.7 90.6 80.8 83.8
OwMatch niu2024owmatch 93.0 95.9 94.4 74.5 55.9 65.1 91.7 72.0 81.8
OwMatch+ niu2024owmatch 96.5 97.1 96.8 80.1 63.9 71.9 91.5 79.6 85.5
ZSCLIP CLIP 88.0 88.1 79.3 57.4 53.1 46.0 89.5 87.9 87.7
CoOp coop 91.1 90.4 81.3 65.4 51.0 48.4 89.7 83.2 85.5
PromptSRC promptsrc 93.4 91.1 83.9 73.0 63.2 59.7 91.7 89.5 89.5
Datasets and Experimental Setup. We conduct evaluations on CIFAR-10/100 cifar , and ImageNet-100 cao2021open . In each dataset, the classes are divided into known (base) and new classes, with the proportion of new classes set to 50% by default. A portion of the samples from the known classes are labeled according to a given label ratio, while the remaining part, together with all the new class data, formed the unlabeled set. This setup requires the model to simultaneously distinguish and classify both base and new classes. Within the context of pretrain-finetuning paradigm, this is commonly known as the Base-to-New setting. For more detailed settings, please refer to Appendix A.4.
@@
w/ LA LA 53.3 $\pm$ 0.44 70.6 $\pm$ 0.21 49.5 $\pm$ 0.40 67.1 $\pm$ 0.78 30.2 $\pm$ 0.44 48.7 $\pm$ 0.89 26.5 $\pm$ 1.31 44.1 $\pm$ 0.42
FixMatch + LA LA 75.3 $\pm$ 2.45 82.0 $\pm$ 0.36 67.0 $\pm$ 2.49 78.0 $\pm$ 0.91 47.3 $\pm$ 0.58 58.6 $\pm$ 0.36 41.4 $\pm$ 0.93 53.4 $\pm$ 0.32
w/ DARP drap 76.6 $\pm$ 0.92 80.8 $\pm$ 0.62 68.2 $\pm$ 0.24 76.7 $\pm$ 1.35 50.4 $\pm$ 0.65 59.0 $\pm$ 0.53 43.6 $\pm$ 1.05 55.3 $\pm$ 0.84
w/ CReST+ wei2021crest 76.7 $\pm$ 1.31 81.1 $\pm$ 0.57 70.9 $\pm$ 1.14 77.9 $\pm$ 0.71 44.0 $\pm$ 0.66 57.1 $\pm$ 1.10 40.6 $\pm$ 0.52 52.3 $\pm$ 0.30
w/ DASO oh2022daso 77.9 $\pm$ 0.88 82.5 $\pm$ 0.60 71.0 $\pm$ 1.68 79.0 $\pm$ 2.23 50.1 $\pm$ 0.75 61.0 $\pm$ 0.72 43.5 $\pm$ 0.81 56.1 $\pm$ 0.72
FixMatch + ACR ACR 81.6 $\pm$ 0.19 84.1 $\pm$ 0.20 77.0 $\pm$ 1.19 82.0 $\pm$ 0.41 51.1 $\pm$ 0.80 61.3 $\pm$ 0.26 46.1 $\pm$ 0.34 56.5 $\pm$ 0.33
FixMatch + CPE CPE 80.7 $\pm$ 0.96 84.4 $\pm$ 0.29 76.8 $\pm$ 0.53 82.3 $\pm$ 0.36 50.6 $\pm$ 0.45 59.8 $\pm$ 0.13 43.6 $\pm$ 0.88 55.8 $\pm$ 0.65
FixMatch + CCL zhou2024continuous 84.5 $\pm$ 0.38 86.2 $\pm$ 0.35 81.5 $\pm$ 0.99 84.0 $\pm$ 0.21 53.5 $\pm$ 0.49 63.5 $\pm$ 0.39 46.8 $\pm$ 0.45 57.5 $\pm$ 0.55
ZSCLIP CLIP 79.30 $\pm$ 0.00 46.00 $\pm$ 0.00
CoOp coop 80.73 $\pm$ 1.90 84.20 $\pm$ 0.50 80.60 $\pm$ 1.79 82.23 $\pm$ 0.75 49.83 $\pm$ 1.04 55.23 $\pm$ 0.50 44.83 $\pm$ 0.45 52.57 $\pm$ 0.12
PromptSRC promptsrc 86.50 $\pm$ 0.37 88.70 $\pm$ 0.08 86.03 $\pm$ 0.48 87.67 $\pm$ 0.42 60.33 $\pm$ 0.17 64.37 $\pm$ 0.34 58.03 $\pm$ 0.46 62.83 $\pm$ 0.48
@@
   w/ DARP drap 66.9 $\pm$ 1.66 75.4 $\pm$ 1.05 59.9 $\pm$ 2.17 72.3 $\pm$ 0.60 30.5 42.5
   w/ CReST+ wei2021crest 56.0 $\pm$ 3.13 71.5 $\pm$ 0.96 55.4 $\pm$ 3.03 68.5 $\pm$ 1.15 32.5 44.7
   w/ DASO oh2022daso 61.3 $\pm$ 1.84 78.4 $\pm$ 0.96 68.5 $\pm$ 1.78 75.3 $\pm$ 0.44 40.9 55.9
   w/ CCL zhou2024continuous 79.1 $\pm$ 0.43 84.8 $\pm$ 1.05 77.1 $\pm$ 0.33 83.1 $\pm$ 0.18 61.5 67.8
ZSCLIP CLIP 97.4 $\pm$ 0.00 46.00 $\pm$ 0.00
CoOp coop 97.77 $\pm$ 0.19 98.13 $\pm$ 0.17 97.73 $\pm$ 0.25 97.83 $\pm$ 0.09 49.83 $\pm$ 1.04 55.23 $\pm$ 0.50
PromptSRC promptsrc 98.27 $\pm$ 0.05 98.47 $\pm$ 0.05 98.13 $\pm$ 0.12 98.30 $\pm$ 0.08 60.33 $\pm$ 0.17 64.37 $\pm$ 0.34
@@
Datasets and Experimental Settings. We compare different methods on four long-tailed datasets: CIFAR-10-LT cifar , CIFAR-100-LT cifar , STL-10-LT stl10 , and ImageNet-127 cossl . For synthetic long-tailed datasets (CIFAR, STL-10), the class-wise labeled/unlabeled sample counts follow an exponential decay controlled by imbalance ratios $\gamma_{l}$ and $\gamma_{u}$ ACR . For ImageNet-127, a naturally long-tailed dataset, we follow zhou2024continuous and use 10% of the training data as labeled. To ensure fair comparison, we adopt the same splits and protocols as SSL methods niu2024owmatch , and report performance of VLMs in a few-shot setting where the number of shots is determined by the smallest labeled class (e.g., 5-shot for CIFAR-10 when $\gamma_{l}=100$ and $N_{1}=500$). Full setup details and shot derivations are available in Appendix A.5.

### v3

@@ 1 Introduction
Deep neural networks have demonstrated performance comparable to that of humans in certain supervised learning tasks, such as image classification [13, 20, 28]. However, these impressive results rely heavily on large amounts of labeled training data. In many real-world applications, obtaining such labeled data often demands substantial human and financial resources [54]. To reduce the dependence on labeled data and thereby cut costs, researchers have proposed Semi-Supervised Learning (SSL)—an effective approach that leverages unlabeled data to enhance model performance.
Recently, the Pretrain-Finetuning paradigm has gained significant attention, as models pre-trained on large-scale, readily available web data exhibit a strong ability to rapidly adapt to downstream tasks through fine-tuning. In image classification, for example, the semantic alignment acquired during pre-training enables models—such as CLIP [43], a pre-trained vision-language model—to perform zero-shot classification without task-specific supervision. Furthermore, building upon the rich knowledge gained during pre-training, various fine-tuning strategies (e.g., prompt tuning) have been proposed to boost task-specific performance with minimal labeled data [24, 25, 56, 57].
@@ 2 Related Work
Semi-Supervised Learning. The goal of Semi-Supervised Learning is to reduce the cost of data annotation by leveraging unlabeled data, under the assumption that both labeled and unlabeled samples belong to a predefined set of known classes. The primary objective is to assign accurate ground-truth labels to unlabeled instances. Existing SSL methods can be broadly categorized into three main approaches: Entropy Minimization [14, 15], Consistency Regularization[30, 38, 46], and Hybrid or Holistic Methods [4, 5]. Specifically, EntMin [14] promotes confident (i.e., low-entropy) predictions on unlabeled samples by introducing an additional loss term that minimizes prediction uncertainty, regardless of class. Pseudo-Labeling[32] assigns pseudo-labels to unlabeled data based on the model’s own predictions throughout training. FixMatch [48] combines consistency regularization and confidence-based filtering by enforcing consistency between weakly and strongly augmented samples, while only using high-confidence pseudo-labels for training, resulting in significant performance gains with limited labeled data.
Pretrain-Finetuning. The recent progress in pre-trained models has garnered significant attention, particularly with the emergence of Large Language Models like GPT [1] and Vision Language Models like CLIP [43]. These models, pre-trained on vast amounts of data, exhibit remarkable adaptability to downstream tasks, leading to the widespread adoption of the pretrain-finetuning paradigm. In image classification tasks, VLMs [2, 43, 55] are pre-trained on large-scale text-image pairs, leveraging shared embeddings of textual and visual representations to enable zero-shot classification. While these pre-trained VLMs have captured broad, generalized representations of images and text, efficiently adapting them to downstream tasks remains a significant challenge. Inspired by the success of prompt tuning in NLP [21, 33], many researchers have proposed to adapt VLMs by learning the prompts in end-to-end training. As the pioneering work, CoOp [57] for the first time introduces learnable prompts to transfer the task-specific knowledge to VLMs. Building on this idea, MaPLe [24] and PromptSRC [25] extend prompt tuning to both visual and textual encoders, enabling joint vision-language adaptation. In Few-shot SSL, we select CoOp and PromptSRC as the primary baselines.
@@ 3.2 Pretrain-Fintuning
Pre-trained models typically involve two stages before downstream evaluation (see Figure 1(b)). Taking CLIP [43] as an example, it employs an image encoder $f$ and a text encoder $g$ to extract feature representations $f(I)$ and $g(T)$ from an input image $I$ and its corresponding text caption $T$, respectively. During pre-training, the model minimizes a contrastive loss over $M$ image-text pairs:
@@
To enable rapid adaptation to downstream tasks, VLMs [56, 57] fine-tune a set of learnable prompts $P=[t_{1},t_{2},\dots,t_{b}]$, while keeping both the image and text encoders frozen. Given class name embeddings $[c_{1},c_{2},\dots,c_{N}]$, each class representation is formed by concatenating $P$ with $c_{i}$, and the prediction is computed as:
@@ 3.3 Few-Shot SSL
  • Standard Semi-Supervised Setting [48]. In this setting, the labels for both subsets belong to the same known class set $\mathcal{Y}^{l}$. For VLMs, which do not rely on unlabeled data for training, the Few-shots’ SSL framework reduces this setting to a standard few-shot scenario, provided the number of labeled samples remains the same.
  • Open-Set Semi-Supervised Setting [41]. In this setting, the unlabeled dataset $\mathcal{D}_{u}$ may include instances from unseen classes, i.e., $\mathcal{Y}^{l}\subset\mathcal{Y}^{u}$ with $\mathcal{Y}^{\text{new}}=\mathcal{Y}^{u}\setminus\mathcal{Y}^{l}$. Since VLMs do not utilize unlabeled data, the open-set setting in Few-shots’ SSL reduces to the semi-supervised setting.
  • Open-World Semi-Supervised Setting [39]. This setting builds upon the open-set setting, aiming to classify both known and new classes. For VLMs, pre-training endows them with zero-shot classification capability, allowing for direct classification of both known and new classes. Therefore, the Open-world semi-supervised setting is analogous to the commonly used Base-to-Novel setting [56] for VLMs.
  • Long-Tailed Semi-Supervised Setting [29]. This setting features a highly imbalanced class distribution, where head classes have many samples and tail classes have few. The imbalance ratio (IR) is defined as the sample count ratio between the largest and smallest classes. Since VLMs use a fixed number of samples per class in few-shot settings, the head and tail classes become consistent due to the same number of samples. Thus, the Few-shot SSL framework reduces this setting to a standard few-shot scenario by matching the number of shots to that of the tail classes, ensuring fair comparison.
@@ 4.1 Baselines Selection
We selected FixMatch [48] (standard semi-supervised learning), DWD [3] (open-set semi-supervised learning), OwMatch [39] (open-world semi-supervised learning), and CCL [58] (long-tailed semi-supervised learning) as representative methods across different settings. For the pretrain–finetuning paradigm, we consider CLIP [43], a pre-trained VLM, along with its prompt tuning variants—CoOp [57], the first to introduce prompt tuning for CLIP, and PromptSRC [25], a state-of-the-art method. Including these methods in our comparison allows for a direct observation of the lower and upper bounds for pre-trained VLMs in semi-supervised settings, leading to a more comprehensive analysis.
@@
UDA [53] 68.78 $\pm$ 0.43 (10% labels)
FixMatch [48] 71.46 $\pm$ 0.52 (10% labels)
ZSCLIP [43] 66.70 $\pm$ 0.00 (0 labels)
CoOp [57] 73.43 $\pm$ 0.12 (2% labels)
PromptSRC [25] 74.03 $\pm$ 0.09 (2% labels)
@@
Datasets and Experimental Setup. We evaluate both SSL and pre-trained VLMs on CIFAR-10/100 [27], STL-10 [9], and ImageNet [10], under varying annotation levels and augmentation strategies. For SSL methods (e.g., UDA [53], MixMatch [5]), we report the standard results as presented in FixMatch [48]. For VLMs, we simulate few-shot learning by training on the same labeled data only, without accessing unlabeled samples. All experiments on VLMs are conducted using the ViT-B/16 backbone and evaluated over three random seeds. Full implementation details are provided in Appendix A.2.
Experiment Results. As shown in Table 1, SSL methods generally perform well on lower-resolution datasets such as CIFAR-10/100 (32 $\times$ 32), surpassing zero-shot CLIP and its fine-tuning variants. This advantage stems from the difficulty pre-trained VLMs face in extracting meaningful features from lower-resolution inputs. In contrast, pre-trained VLMs excels on higher-resolution datasets like STL-10 (96 $\times$ 96), where even zero-shot CLIP surpasses the best-performing SSL baselines. Moreover, as shown in Table 2(a), on the ImageNet dataset, FixMatch requires a significantly larger proportion of labeled data (around 10%) but still lags behind in accuracy. By comparison, VLMs achieve over 73% accuracy using only 2% of labeled samples, with performance improving as more labeled data becomes available. In addition to accuracy, SSL methods typically involve much longer training times. For instance, training SSL models on CIFAR datasets takes several hours [4, 48], whereas prompt tuning approaches for VLMs [24, 25, 57] can be completed in just a few minutes. These observations highlight the complementary strengths of SSL and pretrain-finetuning paradigms, with the latter generally offering a higher performance ceiling and faster training speed.
Analysis. To investigate how input resolution affects pre-trained models’ performance, we evaluate several methods on the STL-10 dataset at varying resolutions (see Table 2(b)). As resolution increases, all VLMs exhibit substantial accuracy gains, highlighting their sensitivity to image resolution. We also experiment with different visual encoder backbones (presented in Appendix B) and find that ViT-based architectures consistently outperform ResNet-50 across all resolutions, indicating that the performance degradation is not due to an intrinsic limitation of Vision Transformer [11]. In fact, it stems from the fixed input size requirement (e.g., 224 $\times$ 224) of VLMs: low-resolution images must be upsampled via interpolation to satisfy this requirement, a process that inevitably introduces distortions and impairs the models’ ability to extract reliable visual features.
@@
Datasets and Experimental Setup. We evaluate both paradigms on standard open-set benchmarks, including CIFAR-10/100 [27], ImageNet-30 [22], and ImageNet-100 [6]. For CIFAR-10/100, we use 100 labeled images per class from CIFAR-10 as the supervised data and the entire CIFAR-100 dataset as the unsupervised data. For ImageNet-30 and ImageNet-100, we adopt a 50-shot setting, using 50 labeled samples per class from the selected in-distribution (ID) classes. It is important to note that the test set in this setting contains only those classes present in the training data. For the pretrain–finetuning paradigm, the experimental setup and VLM performance remain consistent with those described in the semi-supervised setting. Further details can be found in Appendix A.3.
@@
FixMatch [48] 71.5 50.4 49.5 39.6 23.5 20.3 65.8 36.7 34.9
DS3[17] 77.6 45.3 40.2 55.1 23.7 24.0 71.2 32.5 30.8
CGDL [49] 72.3 44.6 39.7 49.3 22.5 23.6 67.3 33.8 31.9
DTC [19] 53.9 39.5 38.3 31.3 22.9 18.3 25.6 20.8 21.3
RankStats [18] 86.6 81.0 82.9 36.4 28.4 23.1 47.3 28.7 40.3
SimCLR [8] 58.3 63.4 51.7 28.6 21.1 22.3 39.5 35.7 36.9
ORCA [7] 88.2 90.4 89.7 66.9 43.0 48.1 89.1 72.1 77.8
NACH [16] 89.5 92.2 91.3 68.7 47.0 52.1 91.0 75.5 79.6
OpenLDN [44] 95.7 95.1 95.4 73.5 46.8 60.1 89.6 68.6 79.1
OpenCon [50] 89.3 91.1 90.4 69.1 47.8 52.7 90.6 80.8 83.8
OwMatch [39] 93.0 95.9 94.4 74.5 55.9 65.1 91.7 72.0 81.8
OwMatch+ [39] 96.5 97.1 96.8 80.1 63.9 71.9 91.5 79.6 85.5
ZSCLIP [43] 88.0 88.1 79.3 57.4 53.1 46.0 89.5 87.9 87.7
CoOp [57] 91.1 90.4 81.3 65.4 51.0 48.4 89.7 83.2 85.5
PromptSRC [25] 93.4 91.1 83.9 73.0 63.2 59.7 91.7 89.5 89.5
Datasets and Experimental Setup. We conduct evaluations on CIFAR-10/100 [27], and ImageNet-100 [6]. In each dataset, the classes are divided into known (base) and new classes, with the proportion of new classes set to 50% by default. A portion of the samples from the known classes are labeled according to a given label ratio, while the remaining part, together with all the new class data, formed the unlabeled set. This setup requires the model to simultaneously distinguish and classify both base and new classes. Within the context of pretrain-finetuning paradigm, this is commonly known as the Base-to-New setting. For more detailed settings, please refer to Appendix A.4.
@@
w/ LA [37] 53.3 $\pm$ 0.44 70.6 $\pm$ 0.21 49.5 $\pm$ 0.40 67.1 $\pm$ 0.78 30.2 $\pm$ 0.44 48.7 $\pm$ 0.89 26.5 $\pm$ 1.31 44.1 $\pm$ 0.42
FixMatch + LA [37] 75.3 $\pm$ 2.45 82.0 $\pm$ 0.36 67.0 $\pm$ 2.49 78.0 $\pm$ 0.91 47.3 $\pm$ 0.58 58.6 $\pm$ 0.36 41.4 $\pm$ 0.93 53.4 $\pm$ 0.32
w/ DARP [26] 76.6 $\pm$ 0.92 80.8 $\pm$ 0.62 68.2 $\pm$ 0.24 76.7 $\pm$ 1.35 50.4 $\pm$ 0.65 59.0 $\pm$ 0.53 43.6 $\pm$ 1.05 55.3 $\pm$ 0.84
w/ CReST+ [51] 76.7 $\pm$ 1.31 81.1 $\pm$ 0.57 70.9 $\pm$ 1.14 77.9 $\pm$ 0.71 44.0 $\pm$ 0.66 57.1 $\pm$ 1.10 40.6 $\pm$ 0.52 52.3 $\pm$ 0.30
w/ DASO [40] 77.9 $\pm$ 0.88 82.5 $\pm$ 0.60 71.0 $\pm$ 1.68 79.0 $\pm$ 2.23 50.1 $\pm$ 0.75 61.0 $\pm$ 0.72 43.5 $\pm$ 0.81 56.1 $\pm$ 0.72
FixMatch + ACR [52] 81.6 $\pm$ 0.19 84.1 $\pm$ 0.20 77.0 $\pm$ 1.19 82.0 $\pm$ 0.41 51.1 $\pm$ 0.80 61.3 $\pm$ 0.26 46.1 $\pm$ 0.34 56.5 $\pm$ 0.33
FixMatch + CPE [35] 80.7 $\pm$ 0.96 84.4 $\pm$ 0.29 76.8 $\pm$ 0.53 82.3 $\pm$ 0.36 50.6 $\pm$ 0.45 59.8 $\pm$ 0.13 43.6 $\pm$ 0.88 55.8 $\pm$ 0.65
FixMatch + CCL [58] 84.5 $\pm$ 0.38 86.2 $\pm$ 0.35 81.5 $\pm$ 0.99 84.0 $\pm$ 0.21 53.5 $\pm$ 0.49 63.5 $\pm$ 0.39 46.8 $\pm$ 0.45 57.5 $\pm$ 0.55
ZSCLIP [43] 79.30 $\pm$ 0.00 46.00 $\pm$ 0.00
CoOp [57] 80.73 $\pm$ 1.90 84.20 $\pm$ 0.50 80.60 $\pm$ 1.79 82.23 $\pm$ 0.75 49.83 $\pm$ 1.04 55.23 $\pm$ 0.50 44.83 $\pm$ 0.45 52.57 $\pm$ 0.12
PromptSRC [25] 86.50 $\pm$ 0.37 88.70 $\pm$ 0.08 86.03 $\pm$ 0.48 87.67 $\pm$ 0.42 60.33 $\pm$ 0.17 64.37 $\pm$ 0.34 58.03 $\pm$ 0.46 62.83 $\pm$ 0.48
@@
   w/ DARP [26] 66.9 $\pm$ 1.66 75.4 $\pm$ 1.05 59.9 $\pm$ 2.17 72.3 $\pm$ 0.60 30.5 42.5
   w/ CReST+ [51] 56.0 $\pm$ 3.13 71.5 $\pm$ 0.96 55.4 $\pm$ 3.03 68.5 $\pm$ 1.15 32.5 44.7
   w/ DASO [40] 61.3 $\pm$ 1.84 78.4 $\pm$ 0.96 68.5 $\pm$ 1.78 75.3 $\pm$ 0.44 40.9 55.9
   w/ CCL [58] 79.1 $\pm$ 0.43 84.8 $\pm$ 1.05 77.1 $\pm$ 0.33 83.1 $\pm$ 0.18 61.5 67.8
ZSCLIP [43] 97.4 $\pm$ 0.00 46.00 $\pm$ 0.00
CoOp [57] 97.77 $\pm$ 0.19 98.13 $\pm$ 0.17 97.73 $\pm$ 0.25 97.83 $\pm$ 0.09 49.83 $\pm$ 1.04 55.23 $\pm$ 0.50
PromptSRC [25] 98.27 $\pm$ 0.05 98.47 $\pm$ 0.05 98.13 $\pm$ 0.12 98.30 $\pm$ 0.08 60.33 $\pm$ 0.17 64.37 $\pm$ 0.34
@@
Datasets and Experimental Settings. We compare different methods on four long-tailed datasets: CIFAR-10-LT [27], CIFAR-100-LT [27], STL-10-LT [9], and ImageNet-127 [12]. For synthetic long-tailed datasets (CIFAR, STL-10), the class-wise labeled/unlabeled sample counts follow an exponential decay controlled by imbalance ratios $\gamma_{l}$ and $\gamma_{u}$ [52]. For ImageNet-127, a naturally long-tailed dataset, we follow [58] and use 10% of the training data as labeled. To ensure fair comparison, we adopt the same splits and protocols as SSL methods [39], and report performance of VLMs in a few-shot setting where the number of shots is determined by the smallest labeled class (e.g., 5-shot for CIFAR-10 when $\gamma_{l}=100$ and $N_{1}=500$). Full setup details and shot derivations are available in Appendix A.5.